A body of mass 6N rest on a plane inclined at 30 degree to horizontal.
a) what force keeps it from sliding down on the plane?
b) what is the coefficient of the friction?
1
Expert's answer
2019-07-09T13:09:53-0400

1) Holds the force of friction.

2)

"F = \mu N""\mu = F/N""\mu = {mg\sin\alpha \over N} = {mg\sin\alpha \over mg\cos\alpha } = \tan \alpha"

"\mu = \tan 30^o"

"\mu = \sqrt{3} / 3 = 0.577"

Answer: 1) friction force; 2) 0.577